# Temporal Entanglement Transitions in the Periodically Driven Ising Chain

### v2

@@ Abstract
Our findings establish TET as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions. At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features. The universal critical behavior persists whenever the above Z2-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve Z2), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [66] for details).
@@
These findings establish TET as fundamental aspects of non-equilibrium Floquet quantum dynamics. Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [78], however the question of robust universality has been an open question. This is what we have addressed in this work, where our TET arise whenever the driven Hamiltonian and initial state respect global $\mathbb{Z}_{2}$, with critical times determined by sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, yielding a nonanalytic reorganization (quantum phase transition) of the EH with Ising-class critical exponents $\nu=1$. TET define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [66]). The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [79, 80]. The requirement of symmetry-preserving initial conditions with critical times selected by sector-weight degeneracies motivates future exploration of how symmetry and entanglement resources jointly control non-equilibrium criticality structures in quantum resource theories [81]. Beyond the TFIM, the criterion (global symmetry, symmetric initial state, sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in other driven systems with discrete symmetries, including higher-spin chains, symmetry-protected topological phases, and lattice gauge theories, with experimental prospects in cold-atom and trapped-ion platforms [82, 83, 84, 75].

### v3

@@ Abstract
Our findings establish TET as a distinct class of non-equilibrium phenomenon that is uniquely detectable through entanglement measures while remaining hidden from conventional observables, revealing that the EH itself can undergo quantum phase transitions. At high frequencies, we find that the EH gains an intrinsic timescale decoupled from the drive, rendering its Floquet-periodic transitions as genuine steady-state features. The universal critical behavior persists whenever the above $\mathbb{Z}_{2}$-symmetry condition and sector-weight equality are satisfied (including across different equilibrium phases that preserve $\mathbb{Z}_{2}$), confirming a fundamentally non-equilibrium character decoupled from the underlying equilibrium phase diagram (see Supplemental Material (SM) [1] for details).
@@
These findings establish TET as fundamental aspects of non-equilibrium Floquet quantum dynamics. Notably, previous work on Floquet symmetry-protected topological system has demonstrated band crossings in entanglement spectrum [61], however the question of robust universality has been an open question. This is what we have addressed in this work, where our TET arise whenever the driven Hamiltonian and initial state respect global $\mathbb{Z}_{2}$, with critical times determined by sector-weight crossings $w_{+}(t_{c}^{(k)})=w_{-}(t_{c}^{(k)})$, yielding a nonanalytic reorganization (quantum phase transition) of the EH with Ising-class critical exponents $\nu=1$. TET define a class of non-equilibrium criticality distinct from Loschmidt echo singularities and magnetization dynamics (see SM [1]). The exact Jordan-Wigner equivalence between the TFIM and Kitaev chain suggests that these transitions should also manifest in driven topological superconductors, opening pathways to entanglement-based probes of Floquet Majorana physics [13, 75]. The requirement of symmetry-preserving initial conditions with critical times selected by sector-weight degeneracies motivates future exploration of how symmetry and entanglement resources jointly control non-equilibrium criticality structures in quantum resource theories [18]. Beyond the TFIM (despite its free-fermion Jordan-Wigner mapping), our symmetry/sector-competition criterion (global symmetry, symmetric initial state, and sector-weight equality $w_{+}(t)=w_{-}(t)$) suggests analogous transitions in interacting $\mathbb{Z}_{2}$-symmetric chains and, more broadly, in other driven systems with discrete symmetries (higher-spin chains, symmetry-protected topological phases, and lattice gauge theories), with experimental prospects in cold-atom and trapped-ion platforms [25, 9, 42, 11].